Bradley Solutions and Alexander Limited are two well-established suppliers of inexpensive tools. Weekend Projects is a national chain of retail outlets that caters to the occasional fixer-upper who would prefer to get the job done fast rather than investing in a well-appointed toolbox. Weekend Projects wants to find a supplier for a particular toolset that promises to be a big seller. Expected annual sales are 100,000 units (D). Weekend's warehouses operate 50 weeks a year. Management collected data on the two suppliers, which are contained in the table below:

   Annual Freight Costs                      
   Shipping Quantity (Q)                      

Supplier 10,000 25,000 50,000 price/unit Annual holding cost/unit lead time
Bradley $30,000 26,000 19,000 $8.15 $1.63 5 wks
Alexander $35,000 28,000 17,000 $8.15 $1.63 4 wks

admin costs

Bradley: $15,000

Alexander: $10,000

a. Which of the two suppliers would provide the lowest annual cost to Weekend Projects? What shipping quantity would you suggest?

Using (Alexander or Bradley ) and a shipping quantity of (25,000, 50,000, or 10,000) units is the lowest cost alternative, with annual total costs to Weekend Projects of $ ______

(Enter your response as an integer.)

b. Before management could make a decision, another option became available. Zelda Tools offered the toolset for only $8.00 however, the lead time is longer than the other two suppliers. Zelda is a new supplier and has not been in the industry very long. Additional data for Zelda are in the table below:

Annual Freight Costs

Shipping Quantity (Q)

Supplier

10,000

25,000

50,000

Price/Unit (p)

Annual Holding Cost/Unit (H)

Lead Time (L) (wks)

Annual Administrative Cost

Zelda

$45,000

$27,000

$19,000

$8.00

$1.60

6

?

Management has begun to assess the administrative costs to manage the contract with Zelda. What is the lowest level of administrative costs at which Weekend Projects would be indifferent between using Zelda versus the option you chose in part (a)?

Weekend Projects would be indifferent if administrative costs are $ _______

(Enter your response as an integer.)
Answer #1

Total annual cost = Freight cost + Purchase cost + Holding cost of Cycle stock + Holding cost of Transit stock + Administrative cost

a)

Total cost of Bradley (for Q=10000) = 30000 + 100000*8.15 + (10000/2 + 100000*5/50)*1.63 + 15000 = $ 884,450

Total cost of Bradley (for Q=25000) = 26000 + 100000*8.15 + (25000/2 + 100000*5/50)*1.63 + 15000 = $ 892,675

Total cost of Bradley (for Q=50000) = 19000 + 100000*8.15 + (50000/2 + 100000*5/50)*1.63 + 15000 = $ 906,050

Total cost of Alexander (for Q=10000) = 35000 + 100000*8.15 + (10000/2 + 100000*4/50)*1.63 + 10000 = $ 886,190

Total cost of Alexander (for Q=25000) = 28000 + 100000*8.15 + (25000/2 + 100000*4/50)*1.63 + 10000 = $ 891,415

Total cost of Alexander (for Q=50000) = 17000 + 100000*8.15 + (50000/2 + 100000*4/50)*1.63 + 10000 = $ 900,790

Of the above, minimum total cost is $ 884,450 of Bradley for Q=10000. Therefore,

Using Bradley, and a shipping quantity of 10000 units is the lowest cost alternative, with annual total costs of $ 884,450

b)

Calculate the total cost (except admin cost) of Zelda for the 3 different order quantities

Total cost of Zelda (for Q=10000) = 45000 + 100000*8.00 + (10000/2 + 100000*6/50)*1.60 = $ 872,200

Total cost of Zelda (for Q=25000) = 27000 + 100000*8.00 + (25000/2 + 100000*6/50)*1.60 = $ 866,200

Total cost of Zelda (for Q=50000) = 19000 + 100000*8.00 + (50000/2 + 100000*6/50)*1.60 = $ 878,200

We see that Total cost of Zelda is the lowest for Q=25000

Difference between the lowest cost option chosen in part (a) and the lowest cost of Zelda = 884450 - 866200 = $ 18,250

The lowest level of administrative cost for which Weekend Projects would be indifferent = 884450 - 872200 = $ 12,250